OpenStudy (anonymous):

The length of a rectangle is 2 more than 3 times the width. If the perimeter is 100 meters, what is the width of the rectangle? 13 meters 12 meters 11 meters

Vocaloid (vocaloid):

let w = width let l = length l = 3w + 2 perimeter = 2(l+w) = 2(3w + 2 + w) = 2(4w + 2) = 100 solve for w

OpenStudy (tkhunny):

So sad now that we will never know if @AlgebraHelp123 can do the interesting part. Most anyone can solve for 'w' after it is all laid out.

OpenStudy (anonymous):

Whats wrong with needing help?

OpenStudy (tkhunny):

Nothing wrong with needing help. Plenty wrong with not getting help and believing that you did.

OpenStudy (anonymous):

Any way, Thanks you @Vocaloid i got it.

OpenStudy (tkhunny):

Before you were given this problem, you should already have gained a background in solving simple equations. This section is intended to teach you how to set up those equations. Someone else did that for you. You have missed the lesson.

Vocaloid (vocaloid):

I see your point, so, @AlgebraHelp123, do you understand what I did?